### v1

@@ 5.2 CP-ALS Performance
We compare our ALTO-based CP-ALS algorithm to an oracle that selects the best performing implementation in the state-of-the-art libraries SPLATT and ParTI for each tensor. The oracle considers two distinct types of tensor formats: 1) mode-agnostic or general formats (COO, HiCOO-SB10, and HiCOO-SB14), which use one tensor copy, and 2) mode-specific formats (CSF and CSF-tile), which keep multiple tensor copies (one per mode) for best performance.
Figure 9 demonstrates that ALTO outperforms the best mode-agnostic as well as mode-specific formats in terms of the speedup of tensor operations (MTTKRP on all modes). In addition, the results indicate that ALTO can effectively avoid synchronization and utilize the larger fast memories on SPR (relative to ICX) to further improve the performance compared to the state-of-the-art libraries. Specifically, compared to the best mode-agnostic formats, ALTO achieves and speedup on the ICX and SPR CPUs, respectively. Although the mode-specific (CSF-based) formats require substantial storage to keep multiple tensor copies, ALTO still delivers and geometric mean speedup on the ICX and SPR CPUs, respectively. Furthermore, ALTO shows scalable performance for the sparse tensors with high data reuse. Compared to the sequential version, it achieves up to and speedup on the 72-core ICX and 104-core SPR CPUs, respectively. For the other tensors (with limited/medium data reuse), ALTO is bounded by the memory bandwidth, and as a result it realizes an average speedup of around and on ICX and SPR, respectively.
@@ 5.3 CP-APR Performance
We compare our ALTO-based CP-APR algorithm to the state-of-the-art SparTen library for non-negative count tensors. SparTen uses a variant of the COO sparse format that keeps indexing as well as scheduling or permutation arrays for every tensor mode, which requires more than double the storage of COO [27]. In addition, SparTen computes CP-APR using the traditional method that stores and then reuses intermediate results rather than recomputing them. As a result, for large sparse tensors, such as reddit, SparTen fails to compute the CP-APR model on the ICX platform. In contrast, ALTO-based CP-APR supports both recomputing (ALTO-OTF) or storing and then reusing (ALTO-PRE) intermediate results, which enables our CP-APR implementation to handle large tensors and high decomposition ranks.
Figures 9(a) and 9(b) show the parallel performance of ALTO-based CP-APR compared to SparTen on the ICX and SPR CPUs, respectively. As the vast majority of time (more than 99 %) is spent in the model update () kernel (see Algorithm 5), the parallel performance is evaluated based on the computation time of this tensor kernel. Like CP-ALS, ALTO-based CP-APR realizes more scalable performance for tensors with high fiber reuse, and it further improves the performance on SPR relative to ICX by leveraging the larger fast memories on SPR while avoiding synchronization when possible. Therefore, ALTO delivers substantial performance gains compared to the SparTen library, achieving and speedup on the ICX and SPR CPUs, respectively. Furthermore, as the tensor size and data reuse increase, our on-the-fly (ALTO-OTF) algorithmic variant not only outperforms the traditional pre-computing approach (ALTO-PRE) but also realizes better scalability, even when the intermediate results can fit in memory.
@@ 5.4 Performance Characterization
Unlike prior compressed tensor formats, the parallel performance of ALTO depends on the inherent data reuse of sparse tensors rather than the spatial distribution of their nonzero elements. To better understand the performance characteristics of the ALTO format, we created a Roofline model [46] for the test platform and collected performance counters across a set of representative parallel runs. For the Roofline model, an upper performance limit is given by , where is the theoretical peak performance, is the peak memory bandwidth, and is the operational intensity (i.e., the ratio of executed floating-point operations per byte). In addition, we enhance our Roofline model to consider the bandwidth of the different cache levels. The L2/L3 cache and main memory bandwidth in our model are measured using likwid-bench. However, since L1 bandwidth measurements are error-prone, we use the theoretical L1 load bandwidth of two cache lines per cycle per core for this particular roofline. The peak performance, , is calculated based on the ability of the cores to execute two fused multiply-add (FMA) instructions on eight-element double precision vector registers (due to the availability of AVX-512) per cycle.
@@
Since prior work already details the performance analysis of the CP-ALS algorithm and its MTTKRP kernel [31], we focus on characterizing the performance of the CP-APR algorithm in this study. Figure 11 shows the performance of the parallel -ALTO kernel (Algorithm 5) for several representative tensors. To quantify the the operational intensity, we calculated the required data movement from/to main memory as for on-the-fly computation (ALTO-OTF) and for pre-computation (ALTO-PRE), where is the average number of inner iterations, is the number of nonzero elements, is the number of modes, and is the decomposition rank. We obtain the number of FLOPs required for the model update () by measuring hardware performance counters using likwid-perfctr. The results indicate that although such memory-intensive computations suffer from low operational intensity, ALTO can still exceed the peak main memory bandwidth by exploiting the inherent data reuse and by efficiently resolving update conflicts in caches.
@@
Specifically, the Roofline plot shows the performance of ALTO-OTF and ALTO-PRE for two tensors with high data reuse (enron and reddit) and two tensor with limited data reuse (lanl and darpa). For enron, we observe that ALTO provides data access in a manner that allows the computation to be handled mainly from L1 and L2 cache. However, as a medium-sized tensor with high data reuse, it does not benefit from pre-computation and shows superior performance for the on-the-fly algorithm. The reddit tensor, despite having good fiber reuse, is highly sparse and it is the largest tensor in our experiments (with 4.6 billion nonzero elements). This increases the memory pressure and effectively leads to more data accesses from slower memory, which reduces the performance gap between ALTO-OTF and ALTO-PRE relative to the enron tensor. While lbnl is extremely sparse and has limited data reuse, it is also the smallest of all tensors in the experiments. This allows ALTO to handle most of the data from the caches and to benefit from on-the-fly computation; however, the performance of lbnl is lower than denser tensors such as enron. Finally, the darpa tensor, even though being similar in size to enron, has very limited fiber reuse (along mode-3). This leads to a significantly lower performance compared to any of the other tensors, yet the hybrid (recursive and output-oriented) tensor traversal of ALTO still captures some data reuse from caches and allows both ALTO-PRE and ALTO-OTF to realize superior performance, exceeding the main memory bandwidth. For darpa we can observe a slightly better performance when using pre-computation (140 MFLOP/s vs 135 MFLOP/s). Hence, the performance analysis indicates that ALTO-PRE is especially relevant for large tensors that additionally show hyper-sparsity and limited data reuse.
@@ 5.5 Memory Storage
Figure 12 details the relative storage of the different sparse tensor representations compared to the COO format. Due to its efficient linearization, as detailed in 3.1, the ALTO format requires less storage than mode-specific (CSF and CSF-tile) and raw (COO) formats for all investigated tensors. The tree-based, mode-specific formats (CSF and CSF-tile) consume significantly more storage space than COO because they require multiple tensor representations for the different mode orientations. While it can be beneficial for computation, imposing a tilling over the tensors (as done by CSF-tile) increases memory storage. The memory consumption of the block-based formats (HiCOO-SB10/SB14) highly depends not only on the spatial distribution of the nonzero elements, but also on the block and superblock sizes. Compared to the COO format, HiCOO can reduce the memory footprint of tensors when the resulting blocks are relatively dense, i.e., the number of nonzero elements per block is high. However, for hyper-sparse tensors such as deli, nell-1, amazon, and reddit, HiCOO requires more storage by up to a factor of .
@@ 5.6 Format Generation Cost
Figure 13 details the generation cost of the different sparse tensor representations from a sparse tensor in the COO format on the SPR platform. Instead of processing nonzero elements in a multi-dimensional form as HiCOO and CSF formats, ALTO works on a linearized representation of tensors, which needs substantially lower number of comparison operations to sort nonzero elements. Furthermore, the HiCOO formats require additional clustering of elements based on their multi-dimensional coordinates, as well as scheduling the blocks and superblocks for avoiding conflicts. Therefore, ALTO achieves substantial geometric mean speedup compared to HiCOO-SB10 (50x), HiCOO-SB14 (72x), CSF-tile (10x), and CSF (6x).
@@ 6 Related Work
Our mode-agnostic ALTO format was motivated by the linearized coordinate (LCO) format [37], which also flattens sparse tensors but in a mode-specific way, i.e., along a given mode orientation. Hence, LCO requires either multiple tensor copies or permuting tensors for efficient computation. Additionally, the authors limit their focus to sequential algorithms, and it is not clear how LCO can be used to efficiently parallelize sparse tensor computations.
Researchers proposed variants [26, 27] of the COO format to reduce the synchronization overhead across threads using mode-specific scheduling/permutation arrays. The state-of-the-art SparTen library [22] employs a COO variant [27] to accelerate the decomposition of non-negative count tensors across different hardware architectures. However, these COO variants adversely affect the input data locality and lead to random access of the nonzero elements [27], especially for sparse tensors with high fiber reuse. Moreover, keeping fine-grained scheduling information for all tensor modes can more than double the memory consumption, compared to the COO format [27].
@@
Alternatively, the ParTI library uses the mode-agnostic, block-based HiCOO format [18] to decompose sparse tensors using only one tensor copy. However, HiCOO is highly sensitive to the characteristics of sparse tensors as well as the block size. Due to the irregular (skewed) data distributions in sparse tensors, the number of nonzero elements per block varies widely across HiCOO blocks, even after expensive mode-specific tensor permutations which in practice further increase workload imbalance [28]. As a result, when tensors are highly sparse, HiCOO can consume more storage than the COO format [18]. Moreover, using small data types for indexing nonzero elements within a block can end up under-utilizing the compute units and memory bandwidth in modern parallel architectures [47, 48], which are optimized for wide memory transactions [49].
SpTFS [19] utilizes machine learning [50, 51] to predict the best of COO, HiCOO, and CSF formats to compute MTTKRP for a given sparse tensor. STeF [52] leverages the mode-specific CSF format to accelerate all-modes MTTKRP by memoization of partial MTTKRP results. Nevertheless, the additional space required for memoization can be more than double the memory storage of the sparse tensor and factor matrices, which limits STeF applicability to small- and medium-scale tensors [52]. FLYCOO [53, 54] extends the COO format to memory-constrained platforms (such as FPGAs) by processing a tensor into small equal-sized shards. However, FLYCOO requires dynamic mode-specific tensor remapping/reordering and it needs more storage than COO to keep sharding information for every mode.
@@ 7 Conclusion
To overcome the limitations of existing compressed tensor formats, this work introduced the ALTO format, a compact mode-agnostic representation to efficiently encode higher-order sparse data of arbitrary dimensionality and irregular distributions. Thanks to their adaptive tensor traversal and superior workload balance and data reuse, our ALTO-based parallel algorithms for decomposing normally distributed data (CP-ALS) and non-negative count data (CP-APR) delivered an order-of-magnitude speedup over the best mode-agnostic formats while requiring of COO storage. Moreover, ALTO achieved geometric mean speedup over the best mode-specific formats with of their tensor storage. Our future work will investigate distributed-memory platforms and other common sparse tensor algorithms, besides tensor decomposition.

### v2

@@ 5.2 CP-ALS Performance
We compare our ALTO-based CP-ALS algorithm to a variety of CP-ALS implementations in the state-of-the-art libraries SPLATT, ParTI, and STeF for each tensor dataset. The implementations can be grouped into three categories: 1) mode-agnostic or general formats (COO, HiCOO-SB10, and HiCOO-SB14), which use one tensor copy, 2) mode-specific formats (CSF and CSF-tile), which keep multiple tensor copies (one per mode) for best performance, and 3) memoization techniques (STeF), which retain intermediate results across modes along with the tensor representation to reduce computations.
Figures 9 and 10 show that ALTO outperforms the best mode-agnostic and mode-specific formats as well as memoization schemes in terms of the speedup of tensor operations (MTTKRP on all modes). In addition, the results indicate that ALTO can effectively reduce synchronization and utilize the larger caches on SPR (relative to ICX) to further improve the performance compared to the state-of-the-art libraries. Specifically, ALTO achieves and geometric mean (GEOMEAN) speedup on the ICX and SPR CPUs, respectively, compared to the best mode-agnostic formats. Although the mode-specific (CSF-based) formats require substantial storage to keep multiple tensor copies, ALTO still delivers and geometric mean speedup on the ICX and SPR servers, respectively. While memoization methods reduce computations, it comes at the cost of increasing memory traffic and consuming substantial amount of extra memory, which significantly limit their scalability. As a result, ALTO realizes and geometric mean speedup over STeF on the ICX and SPR CPUs, respectively, while effectively handling all large-scale tensors that cause out-of-memory errors with STeF. Furthermore, ALTO shows scalable performance for the sparse tensors with high data reuse. Compared to its sequential version, ALTO achieves up to and speedup on the 72-core ICX and 104-core SPR CPUs, respectively. For the other tensors (with limited/medium data reuse), ALTO is bounded by memory bandwidth, and as a result it has an average speedup of around and on ICX and SPR, respectively. Conversely, the performance of the previous approaches is highly variable across data sets as it depends on the shape of sparse tensors as well as the spatial distribution of their nonzero elements rather than their inherent data reuse [30]. Specifically, the tree-based (CSF, CSF-tile, and STeF) and block-based (HiCOO-SB10/SB14) methods depend on grouping the nonzero elements into tensor slices and blocks for effective compression. As illustrated in Figure 1, finding balanced clusters of nonzero elements in sparse tensors is highly unlikely. Thus, the prior tree- and block-based techniques suffer from workload imbalance and inefficient compression, which in turn lead to limited parallel performance when scaling to a large number of cores.
@@ 5.3 CP-APR Performance
We compare our ALTO-based CP-APR algorithm to the state-of-the-art SparTen library for non-negative count tensors. SparTen uses a variant of the COO format that keeps indexing as well as scheduling arrays for every tensor mode, which requires more than double the storage of COO [26]. In addition, SparTen computes CP-APR using the traditional method that stores and then reuses intermediate results rather than recomputing them. Hence, for large sparse tensors, such as reddit, SparTen fails to compute the CP-APR model on the ICX platform, even with 256 GiB of memory. In contrast, ALTO supports both recomputing (ALTO-OTF) or storing and then reusing (ALTO-PRE) intermediate results, which enables our CP-APR implementation to handle large-scale tensors.
Figure 11 shows the parallel performance of ALTO-based CP-APR compared to SparTen on the ICX and SPR CPUs. As the vast majority of time (more than 99 %) is spent in the model update () kernel (see Algorithm 5), the performance is evaluated based on the computation time of this tensor kernel. Note that ALTO-PRE and ALTO-OTF in Figure 11 represent the speedup achieved when the respective algorithms are used for all input tensors, and ALTO represents the speedup achieved when our adaptive memory management heuristic (4.3) is used to choose between the two algorithms to maximize performance. Like CP-ALS, ALTO-based CP-APR realizes more scalable performance for tensors with high fiber reuse, and it further improves the performance on SPR relative to ICX by reducing synchronization and leveraging the larger fast memories on SPR. Therefore, ALTO delivers substantial performance gains compared to the SparTen library, achieving and speedup on the ICX and SPR CPUs, respectively. Furthermore, as the tensor size and data reuse increase, our on-the-fly (ALTO-OTF) algorithmic variant not only outperforms the traditional pre-computing approach (ALTO-PRE) but also realizes better scalability, even when the intermediate results can fit in memory.
@@ 5.4 Performance Characterization
Unlike prior compressed tensor formats, the parallel performance of ALTO depends on the inherent data reuse of sparse tensors rather than the spatial distribution of their nonzero elements. To better understand the performance characteristics of ALTO, we created a Roofline model [46] for the SPR platform and collected performance counters across a set of representative parallel runs. For the Roofline model, an upper performance limit is given by , where is the peak performance, is the peak memory bandwidth, and is the operational intensity (i.e., the ratio of floating-point operations per byte). Moreover, we enhance our Roofline model to consider the cache bandwidth. The L2/L3 cache and main memory bandwidth are measured using likwid-bench from the Likwid tool suite. Since L1 bandwidth measurements are error-prone, we use the theoretical L1 bandwidth of two cache lines per cycle per core. The peak performance, , is calculated based on the ability of the cores to execute two fused multiply-add (FMA) instructions on eight-element double precision vector registers (due to the availability of AVX-512) per cycle.
@@
Since prior work detailed the performance analysis of the CP-ALS algorithm and its MTTKRP kernel [30], we focus on characterizing the performance of CP-APR in this study. Figure 12 shows the performance of the parallel -ALTO kernel (Algorithm 5) for several representative tensors. To quantify the operational intensity, we calculated the required data movement from/to main memory as for on-the-fly computation (ALTO-OTF) and for pre-computation (ALTO-PRE), where is the average number of inner iterations, is the number of nonzero elements, is the number of modes, and is the decomposition rank. We obtain the number of FLOPs required for the model update () by measuring hardware performance counters using likwid-perfctr from the Likwid tool suite. The results indicate that although such memory-intensive computations suffer from low operational intensity, ALTO can still exceed the peak main memory bandwidth by exploiting the inherent data reuse and by efficiently resolving update conflicts in caches.
@@
Specifically, the Roofline plot shows the performance of ALTO-OTF and ALTO-PRE for two tensors with high data reuse (enron and reddit) and two tensor with limited data reuse (lanl and darpa). For enron, we observe that ALTO provides data access in a manner that allows the computation to be handled mainly from L1 and L2 cache. However, as a medium-sized tensor with high data reuse, it does not benefit from pre-computation and shows superior performance for the on-the-fly algorithm. The reddit tensor, despite having good fiber reuse, is highly sparse and it is the largest tensor in our set of experiments (with 4.6 billion nonzero elements). This increases the memory pressure and effectively leads to more data accesses from slower memory, which reduces the performance gap between ALTO-OTF and ALTO-PRE relative to the enron tensor.
While lbnl is extremely sparse and has limited data reuse, it is also the smallest of all tensors in the experiments. This allows ALTO to handle most of the data from the caches and to benefit from on-the-fly computation; however, the performance of lbnl is lower than denser tensors such as enron. Finally, the darpa tensor, even though being similar in size to enron, has very limited fiber reuse (along mode-3). This leads to a significantly lower performance compared to any of the other tensors, yet the hybrid (recursive and output-oriented) tensor traversal of ALTO still captures some data reuse from caches and allows both ALTO-PRE and ALTO-OTF to realize superior performance, exceeding the main memory bandwidth. For darpa we can observe a slightly better performance when using pre-computation. Hence, the performance analysis indicates that ALTO-PRE is especially relevant for large tensors that additionally show hyper-sparsity and limited data reuse.
@@ 5.5 Memory Storage
Figure 13 details the relative storage of the different sparse tensor formats compared to COO. Due to its efficient linearization, as detailed in 3.1, ALTO requires less storage than the CSF, CSF-tile and raw (COO) formats for all investigated tensors. The tree-based, mode-specific formats (CSF and CSF-tile) consume significantly more storage space than COO because they require multiple tensor representations for the different mode orientations. While it can be beneficial for computation, imposing a tilling over the tensors (as done by CSF-tile) increases memory storage. The memory consumption of the block-based formats (HiCOO-SB10/SB14) highly depends not only on the spatial distribution of the nonzero elements, but also on the block and superblock sizes. Compared to the COO format, HiCOO can reduce the memory footprint of tensors when the resulting blocks are relatively dense, i.e., the number of nonzero elements per block is high. However, for hyper-sparse tensors such as deli, nell-1, amazon, and reddit, HiCOO requires more storage by up to a factor of . While the tensor format of STeF (“STeF (tensor)” in Figure 13) only requires storage on-par or even smaller than ALTO, the additional memory needed for memoization leads to a higher memory footprint (“STeF”) compared to ALTO in all cases.
@@ 5.6 Format Generation Cost
Figure 14 details the generation cost of the different sparse tensor representations from a sparse tensor in the COO format on the SPR platform. Instead of processing nonzero elements in a multi-dimensional form as HiCOO and CSF, ALTO works on a linearized representation that needs substantially lower number of comparison operations to sort nonzero elements. Furthermore, the HiCOO formats require additional clustering of elements based on their multi-dimensional coordinates, as well as scheduling of the blocks and superblocks for avoiding conflicts, while STeF requires additional sorting of the nonzero elements along a specific mode order for best performance. Thus, ALTO achieves substantial geometric mean speedup for format generation compared to HiCOO-SB10 (), HiCOO-SB14 (), CSF-tile (), CSF (), and STeF ().
@@ 6 Related Work
Our mode-agnostic ALTO format was motivated by the linearized coordinate (LCO) format [39], which also flattens sparse tensors but in a mode-specific way, i.e., along a given mode orientation. Hence, LCO requires either multiple tensor copies or permuting tensors for efficient computation. Additionally, the authors limit their focus to sequential algorithms, and it is not clear how LCO can be used to efficiently execute parallel sparse tensor computations.
Researchers proposed variants [21, 26] of the COO format to reduce the synchronization overhead using mode-specific scheduling arrays. The state-of-the-art SparTen library [22] uses a COO variant [26] to accelerate the decomposition of non-negative count tensors across different hardware architectures. However, these COO variants adversely affect the input data locality and lead to random access of the nonzero elements [26], especially for sparse tensors with high data reuse. Moreover, keeping fine-grained scheduling information for all tensor modes can more than double the memory consumption, compared to the COO format [26].
@@
Alternatively, the ParTI library uses the mode-agnostic, block-based HiCOO format [15] to decompose sparse tensors using only one tensor copy. Yet, HiCOO is highly sensitive to the characteristics of sparse tensors as well as the block size. Due to the irregular (skewed) data distributions in sparse tensors, the number of nonzero elements per block varies widely across HiCOO blocks, even after expensive mode-specific tensor permutations which in practice further increase workload imbalance [27]. As a result, when tensors are highly sparse, HiCOO can consume more storage than the COO format [15]. Moreover, using small data types for indexing nonzero elements within a block can end up under-utilizing the compute units and memory bandwidth in modern parallel architectures [47, 48], which are optimized for wide memory transactions [49].
STeF [38] leverages the mode-specific CSF format to accelerate all-modes MTTKRP by memoization of partial MTTKRP results. Nevertheless, the additional space required for memoization can be more than double the memory storage of the sparse tensor and factor matrices, which limits STeF applicability to small- and medium-scale tensors [38]. SpTFS [16] utilizes machine learning [50, 51] to predict the best of COO, HiCOO, and CSF formats to compute MTTKRP for a given sparse tensor. FLYCOO [52, 53] extends the COO format to memory-constrained platforms (such as FPGAs) by processing a tensor into small equal-sized shards. However, FLYCOO requires dynamic mode-specific tensor remapping/reordering and it needs more storage than COO to keep sharding information for every mode.
@@ 7 Conclusion
To overcome the limitations of existing sparse tensor formats, this work introduced ALTO, a compact mode-agnostic format to efficiently encode higher-order tensors of irregular shapes and data distributions. Thanks to their adaptive tensor traversal and superior workload balance and data reuse, our ALTO-based parallel algorithms for decomposing normally distributed data (CP-ALS) and non-negative count data (CP-APR) delivered an order-of-magnitude speedup over the best mode-agnostic formats while requiring of COO storage. Moreover, ALTO achieved and geometric mean speedup over the best mode-specific and memoization methods, respectively, while needing between to of their overall storage. Our future work will investigate distributed-memory platforms and other common sparse tensor algorithms, besides tensor decomposition.